cosx dy/dx=y 2 (1-sinx) cosx

cosx dy/dx=y(1-sinx) cosx

Grade:

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

cosx.\frac{dy}{dx} = y^{2}.(1-sinx)cosx
\frac{dy}{dx} = y^{2}.(1-sinx)
\frac{dy}{y^{2}} = (1-sinx)dx
\int \frac{dy}{y^{2}} = \int (1-sinx)dx
\frac{-1}{y} = x + cosx + constant
y = \frac{-1}{x + cosx + constant}
Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

cosx.\frac{dy}{dx} = y^{2}(1-sinx)cosx
\frac{dy}{dx} = y^{2}(1-sinx)
\frac{dy}{y^{2}} = (1-sinx)dx
\int \frac{dy}{y^{2}} = \int (1-sinx)dx
\frac{-1}{y} = x+cosx+constant
y = \frac{-1}{x+cosx+constant}